A company is planning to replace old equipment with new, more efficient equipment. The company spent $120,000 on a market study and consulting a few months ago. The new equipment will cost $2,200,000. In order to make the equipment operable, the company must pay $60,000 for installation and $90,000 for necessary training. With the anticipated growth due to the equipment replacement, the company has to invest $110,000 in working capital. The equipment will be depreciated via the straight-line depreciation method down to $400,000 over 20 years. What is the annual depreciation expense?

A)$97,500

B)$90,000

C)$117,500

D)$103,000

E)$109,000
Answer #1

total cost = (2,200,000 + 60,000 + 90,000) = 2,350,000

so annual depreciation expense = (2,350,000 - 400,000)/20

annual depreciation expense = 97,500